OpenStudy (anonymous):

Guys!I need your help and answers!Thanks! Lori bought a 2-liter bottle of soda. Yesterday she drank 72 centiliters of the soda. How much of the soda was left? a.1 liter 1 deciliter 4 centiliters b.2 liters 2 deciliters 8 centiliters c.1 liter 7 deciliters 2 centiliters d.1 liter 2 deciliters 8 centiliters

jimthompson5910 (jim_thompson5910):

There are 100 centiliters in a liter (centi means 100, think cent as in penny)

jimthompson5910 (jim_thompson5910):

So 1 liter = 100 centiliters 2*1 liter = 2*100 centiliters 2 liters = 200 centiliters

jimthompson5910 (jim_thompson5910):

If you start off with 2 liters or 200 centiliters, and you drink 72 centiliters, how many centiliters are left?

OpenStudy (anonymous):

128..

jimthompson5910 (jim_thompson5910):

128 centiliters, good

jimthompson5910 (jim_thompson5910):

now we must break that down to get the answer in the form they want

jimthompson5910 (jim_thompson5910):

How many full liters are in 128 centiliters?

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

subtract off 100 from 128 to get 128 - 100 = 28 So the remaining portion that doesn't go with the full liter (100 centiliters) is 28 centiliters In other words, we have 28 centiliters left over

jimthompson5910 (jim_thompson5910):

how many full deciliters are in 28 centiliters ?

jimthompson5910 (jim_thompson5910):

if you're not sure, then 10 deciliters = 1 liter (deci means tenth) 10 deciliters = 100 centiliters ... replace 1 liter with 100 centiliters 1 deciliter = 100/10 centiliters ... divide both sides by 10 1 deciliter = 10 centiliters In short, 1 deciliter = 10 centiliters